OpenStudy (anonymous):

The base of a triangle is 8cm greater than the height (h+8) the area is 42cm^2. find the height and length of the base.

OpenStudy (anonymous):

im trying to use a=1/2*b*h and am getting 38 for the height which is all wrong.

OpenStudy (lalaly):

area of triangle = 1/2 * base *height 42= 1/2 * (h+8) h 84=h^2 + 8h h^2+8h-84=0 (h-6)(h+14)=0 h=6 or -14 the height can never be negative... so its 6

OpenStudy (lalaly):

base= 6+8 = 14

OpenStudy (anonymous):

thank you